OpenStudy (anonymous):

find (f)(g) if f(x) = 3x and g(x) = x + 1

OpenStudy (studygurl14):

(3x)(x + 1) \[3x^2+3x\]

OpenStudy (jdoe0001):

\(\bf f(x) = 3x \qquad g(x) = x + 1\qquad (f)(g) \implies (3x)(x + 1)\)

OpenStudy (studygurl14):

That's if you meant f(x) times g(x)

OpenStudy (anonymous):

the options are 3x + 3 3x + 1 4x + 1 3x2 + 1

OpenStudy (studygurl14):

oh, then your question must have been f of g, which is f(g(x)) Which is: 3(x+1) = 3x + 3

OpenStudy (anonymous):

so its... 3x+1?

geerky42 (geerky42):

"3(x+1) = 3x + 3" So how did you get 3x+1?

OpenStudy (anonymous):

its the closest lol, im just guessing. i dont understand this

OpenStudy (studygurl14):

It is 3x + 3. That's answer choice #1